• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Singh

    Log entry time 06:37:49 on August20,2003

    Entry number 112039

    This entry is a followup to: 112034

    keyword=carbon target problem? pt 2

    here are the ph_tg and th_tg spectra from the same two runs mentioned before. it almost seems as if ph_tg and th_tg are swapped for run 23751 (after recovery). wierd. the magic 8 says "it is decidecly so."

    in addition we noticed that the T3 rates for the two runs are quite different. before the crash (run# < 23749), the T3 rates were about 640 Hz at a beam current of about 3 microA. after the recovery (run# > 23751), the T3 rates have been about 70 Hz at a beam current about 4.7 microA. the magic 8 says "outlook not so good."

    note that we did not change the kinematics at at all. for both runs, the beam energy is 2235 MeV, and the spec setting is 2170 Mev / 18 deg.


    FIGURE 1

    FIGURE 2